OpenStudy (anonymous):

how to solve this equation? 4x^(1/3)=2x^(-2/3)

OpenStudy (anonymous):

Solve for x over the real numbers: 4 x^(1/3) = 2/x^(2/3) Move everything to the left hand side. Subtract 2/x^(2/3) from both sides: 4 x^(1/3)-2/x^(2/3) = 0 Simplify 4 x^(1/3)-2/x^(2/3) = 0 by making a substitution. Simplify and substitute y = x^(1/3): 4 x^(1/3)-2/x^(2/3) = 4 x^(1/3)-(2)/(x^(1/3))^2 = 4 y-2/y^2 = 0: 4 y-2/y^2 = 0 Write the left hand side as a single fraction. Bring 4 y-2/y^2 together using the common denominator y^2: (2 (2 y^3-1))/y^2 = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 2: (2 y^3-1)/y^2 = 0 Multiply both sides by a polynomial to clear fractions. Multiply both sides by y^2: 2 y^3-1 = 0 Isolate terms with y to the left hand side. Add 1 to both sides: 2 y^3 = 1 Divide both sides by a constant to simplify the equation. Divide both sides by 2: y^3 = 1/2 Elminate the power on the left hand side. Take cube roots of both sides: y = 1/2^(1/3) Perform back substitution on y = 1/2^(1/3). Substitute back for y = x^(1/3): x^(1/3) = 1/2^(1/3) Eliminate the cube root on the left hand side. Raise both sides to the power of three: Answer: | | x = 1/2

OpenStudy (anonymous):

I hope that's something correct.

OpenStudy (anonymous):

sorry, i don;t get it. it looks too messy ` could you please use Equation?

OpenStudy (anonymous):

Is x = 1/2 even remotely the answer?

OpenStudy (anonymous):

yes the answer is 1/2, but i need detail,

OpenStudy (anonymous):

could you please use Equation?

OpenStudy (anonymous):

Solve for x over the real numbers: 4 x^(1/3) = 2/x^(2/3) Subtract 2/x^(2/3) from both sides: 4 x^(1/3)-2/x^(2/3) = 0 Simplify and substitute y = x^(1/3): 4 x^(1/3)-2/x^(2/3) = 4 x^(1/3)-(2)/(x^(1/3))^2 = 4 y-2/y^2 = 0: 4 y-2/y^2 = 0 Bring 4 y-2/y^2 together using the common denominator y^2: (2 (2 y^3-1))/y^2 = 0 Divide both sides by 2: (2 y^3-1)/y^2 = 0 Multiply both sides by y^2: 2 y^3-1 = 0 Add 1 to both sides: 2 y^3 = 1 Divide both sides by 2: y^3 = 1/2 Take cube roots of both sides: y = 1/2^(1/3) Substitute back for y = x^(1/3): x^(1/3) = 1/2^(1/3) Raise both sides to the power of three: Answer: | | x = 1/2

OpenStudy (anonymous):

x^(1/3): 4 x^(1/3)-2/x^(2/3) = 4 x^(1/3)-(2)/(x^(1/3))^2 = 4 y-2/y^2 = 0:

OpenStudy (anonymous):

dont get that part

OpenStudy (mathstudent55):

Here's a simpler way: